OpenStudy (loser66):

Let A, B be sets f:A-->B let A1,A2,....An be subsets of A Give an example show \(f(\cap_i A_i)< \cap_i f( A_i)\) Please help

OpenStudy (loser66):

@myininaya

myininaya (myininaya):

what does that subscript i mean next to the intersection i understand the subscript on the A

OpenStudy (loser66):

that just show A_i is arbitrary

myininaya (myininaya):

\[f(A_1 \cap A_2 \cap \cdots \cap A_n)=A_1 \cap A_2 \cdots \cap A_n \] this is what we want to give an example for ?

OpenStudy (loser66):

and f is a function

myininaya (myininaya):

f:A->B be defined by f(x)=x should work , right?

OpenStudy (loser66):

nope, f (intersection subsets) < intersection of f ( subsets) strictly lesser than

myininaya (myininaya):

omg i didn't see the inequality

OpenStudy (ikram002p):

do u mean order of f(union ) < order (union )

OpenStudy (ikram002p):

typo ,intersection :P

OpenStudy (loser66):

for example A1 = [0,1,2] A2=[0,1] then \(A_1\cap A_2= [0,1]\) let define f(x) =x^2 , then \(f(A_1\cap A_2)\) = [0,1] \(f(A_1) =[0,1,4]\) \(f(A_2)= [0,1] \) and \(f(A_1)\cap f(A_2)=[0,1]\) I always see they are = , not <

OpenStudy (loser66):

I mean \(f(A_1\cap A_2)=f(A_1)\cap f(A_2)=[0,1]\) but my question is : is there any function which makes the left hand side strictly < the right hand side?

myininaya (myininaya):

in that last part you didn't mean to say f(A1) instersect f(A2) right? I don't see any f's on the right hand side or our inequality.

OpenStudy (ikram002p):

hmm i was looking for my note , i thought somewhere this is written as a theorem . \(f(A_i\cap A_j)=f(A_i)\cap f(A_j) \) thats why i asked what does th < notations even mean :O like order or comparing sets hmmm

myininaya (myininaya):

of* not or

OpenStudy (ikram002p):

and see no f in the right side , however this is theorem \(f(\cap_i A_i)=\cap_i (f A_i )\) and what ur trying to show is different :3

OpenStudy (ikram002p):

oh ok :O A1 = [0,2,4] A2=[2,1,3] then A1∩A2=[2] let define f(x) =x! , then f(A1∩A2) = [2] f(A1)=[1,2,24] f(A2)=[2,1,6] f(A1)∩f(A2)=[2,1] does this mean f(A1∩A2)< f(A1)∩f(A2) ?

OpenStudy (loser66):

Perfect!! thanks so much.

OpenStudy (loser66):

but f is not function?

OpenStudy (loser66):

since f(0) = f(1), right?

OpenStudy (ikram002p):

:O yeah discreat xD

myininaya (myininaya):

it is a discrete function (loser , are you think about injective because this isn't an injective function)

OpenStudy (loser66):

Yes, It is good enough to show the case. Since the prof didn't ask for a special function, we can choose any of them to show.

OpenStudy (ikram002p):

ok then it make sense now if f in injective then \(f(\cap_i A_i)=\cap_i (f A_i )\) if u take ur example f=x^2 u could only change sets to make an example for ur question :) ( assuming its not asking for 1-1 or onto )

OpenStudy (loser66):

yes, my bad. hehehe.

OpenStudy (ikram002p):

hehe <3

OpenStudy (loser66):

I need watch a lot of video tapes to master the concept. I can't be failed. hehehe

OpenStudy (ikram002p):

whats ur class ?

OpenStudy (loser66):

real analysis

OpenStudy (ikram002p):

cool ^_^ ok then i would like to see all of ur post even if i dint know answers i would like to review it , ok ? i have took it 3 years ago , so i dont wanna forget it :O

OpenStudy (loser66):

good student.!!